# Tractable Probabilistic Models for Investment Planning

### v1

@@ Abstract
This framework enables direct embedding of chance-constrained optimization into investment planning, enforcing safety or reliability with prescribed confidence levels. TPMs allow both scenario analysis and volatility quantification by compactly representing high-dimensional uncertainties. We demonstrate the approach’s effectiveness through a representative power system planning case study, illustrating computational and reliability advantages over traditional scenario-based models.
@@ 1 Introduction
With the increasing and more tangible effects of climate change, policymakers are intensifying demands for cleaner power systems, compelling grid operators to accelerate the integration of renewable energy into their networks. While renewables provide a feasible pathway to decarbonization and meeting greenhouse gas reduction targets, their weather-dependent and stochastic nature introduces profound operational challenges. Reliability is paramount: supply interruptions can propagate through the economy, disrupt critical infrastructure, and undermine social welfare. At the same time, the variability of renewable sources increases the need for additional reserves, flexibility options, and grid reinforcements, all of which carry significant costs that grow with penetration levels [edenhofer2011renewable, impram2020challenges]. To maintain cost-effectiveness and resilience, system operators must ensure electricity remains affordable while safeguarding against blackouts and extreme events. These dual requirements have made rigorous planning indispensable, with infrastructure investments carefully evaluated to minimize the risk of decisions that could compromise long-term security and resilience of supply.
@@
Planning in power systems has long focused on how much new capacity to build, when to deploy it, and which technologies are most suitable to meet projected growth in demand. Traditionally, modelling these generation and transmission expansion problems relies on deterministic assumptions, where long-term forecasts of demand, fuel prices, and technology costs are fixed inputs [taylor2015convex, kirschen2018fundamentals]. Over time, however, it became evident that such assumptions overlook the fundamental uncertainty associated with economic growth, policy developments, the impacts of climate change, and demand response [paragElectricityMarketDesign2016a]. Consequently, more recent approaches emphasize incorporating uncertainty directly into planning frameworks to ensure that expansion strategies remain viable under a wide range of possible futures.
@@
To address these challenges, planning methodologies have evolved towards stochastic programming and robust optimization [ben2009robust, conejo2010decision, pflug2014multistage, hadjidimitriou2021mathematical], which provide systematic ways of evaluating investment decisions under uncertain demand growth and variable renewable generation [zakaria2020uncertainty, 10931726]. In parallel, structural changes in the sector are shifting attention from large-scale, centralized capacity additions toward the integration of distributed energy resources (DERs), such as rooftop solar, storage systems, and demand response, which require more decentralized and flexible planning approaches [wang2013stochastic, asensio2016joint, MOTTA2024401]. Together, these developments reflect a broader transition in power system planning, where the traditional focus on least-cost expansion now also focuses on the need to ensure system resilience and long-term reliability in the face of uncertain and rapidly changing operating conditions.
@@
In this work, we address the challenges of long-horizon investment planning in power utilities, such as generation and transmission expansion planning, where operators must make decisions decades in advance under profound uncertainty. Forecasting the future energy mix and consumption patterns over such horizons is inherently complex, and current practice often relies on a limited set of discrete scenarios that represent key trajectories of exogenous variables. Statistically, this corresponds to a mixture of Dirac measures, which allows planners to capture a handful of plausible futures but offers little insight into the variability and volatility within each scenario. We propose replacing this paradigm with inference over TPMs integrated into the optimization problem. By compactly representing distributions over high-dimensional uncertainties, TPMs enable reasoning about both scenarios and the volatility around them, and the inference is representable in mixed-integer optimization. We illustrate this approach with a representative case study, demonstrating its feasibility for expansion planning under uncertainty and highlighting its potential for broader application across diverse planning contexts in the energy sector.
@@ 2 Background
Planning in Power Systems. Planning is a cornerstone of power system operation and development, ensuring that electricity supply can meet future demand reliably and at reasonable cost. Because power infrastructure has long lifetimes and requires significant capital investment, planning decisions must be made years or even decades in advance, often under deep uncertainty about technology, policy, and demand evolution. This task has traditionally been formulated as a capacity expansion problem, typically divided into generation expansion planning [DeJonghe2012830, Oderinwale2020] (GEP), transmission expansion planning [Xie2020, Lschenbrand2021] (TEP), and distribution expansion planning [Anjo2018, Gbadamosi2020] (DEP). Each has its own time horizons and decision variables. Collectively, these planning activities form the backbone of long-term power system development, linking near-term operational security with long-term goals such as decarbonization, resilience, and affordability [MOTTA2024401]. Traditional formulations often adopt deterministic structures that model uncertainty under a single problem containing a set of scenarios, which are limited by computational power, leading to commonly picking the most representative cases. As long-term planning increases the uncertainty, it is typical to make some considerations, like focusing only on one of the expansion problems, ignoring the unit commitment constraints, or modelling the transmission system with a DC Optimal Power Flow (DCOPF) approach. The DCOPF is a linearized approximation of the AC power flow equations that simplifies the AC power flow equations, enabling scalable optimization for large planning problems, albeit at the expense of detailed network accuracy [doi:10.1049/iet-gtd.2015.0561]. More recent work emphasizes the need to explicitly capture uncertainty in demand and external drivers such as climate change. For example, the work in [LI2016401] models discrete climate scenarios within a stochastic generation expansion framework to evaluate viable investment strategies across multiple futures.
@@
Together, $X$ and $\Xi$ incorporate uncertainty into the investment decision process to make a final decision aligned with the operator’s primary objective: minimizing costs while keeping the system’s reliability. However, as the space of possible values for each $\xi_{t}$ is extremely large, it is prohibitive to explore them all; thus, it is common to build an estimate ($\hat{\Xi}$) with some representative cases and optimize $x$ over it, yielding to solving one constrained optimization problem for each $\xi\sim\hat{\Xi}$ as follows:
@@
to posteriorly pick the least expensive solution among the selected scenarios. The constraints in this (3) represent the different considerations that restrict the minimization of the primary objective (investment cost) while ensuring respect for all scenario-dependent operational and policy constraints. This compact notation allows the formulation to encompass various system constraints without explicitly enumerating each one.
@@
Relevant Work on Scenario Expansion Planning. In power system planning, stochastic optimization and robust methods represent two complementary pathways to deal with uncertainty. For instance, Giannelos et al. [giannelos2025stochastic] propose a multi-stage stochastic network expansion model that treats both exogenous and endogenous uncertainties, capturing how investments adapt over time under varying conditions. On the robust side, Chen et al. [chenRobustOptimizationTransmission2014] introduced minimax cost and regret metrics to guard against worst-case outcomes. In contrast, Roldán et al. [roldanRobustTransmissionNetwork2019] extended that line by modelling correlated uncertainty sets, making the robust expansion plans more realistic. Other works refine robustness further: Li et al. [Li2016] bring in ramping uncertainty and construction delays, and Velloso et al. [velloso2018distributionally] move toward distributionally robust formulations that sit between full stochastic and worst-case models. More recently, Brenner et al. [brenner2024learning] combine stochastic planning with learning-based approximations to handle computational complexity in large-scale systems. In parallel, Gholami et al. [gholami2022dynamic] explore joint expansion of power and gas infrastructures under uncertainty, illustrating how interdependent systems can be co-optimized.
@@
Chance Constraint Optimization. A central difficulty in power system planning is that these decisions happen under uncertainty. Traditional deterministic formulations provided valuable insights but offered limited protection against the wide range of plausible futures. As the penetration of renewables grows and climate-related risks become more pronounced [paragElectricityMarketDesign2016a], planning frameworks increasingly require explicit methods to incorporate uncertainty and evaluate the robustness of proposed investment strategies. One promising direction is including chance constraints, which consist of enforcing constraints probabilistically rather than for all possible outcomes, an idea previously explored for operation optimization problems.
@@
Chance-constrained formulations have already existed to enhance short-term scheduling. For example, Tan et al. [tan2020dual] developed a dual-timescale generation scheduling framework that integrates wind power, demand response, and energy storage, using chance constraints to procure sufficient ramping capacity under uncertainty. Similarly, Wang et al [wang2013stochastic] studied stochastic unit commitment with uncertain demand response, where a random variable models customer price elasticity and a chance constraint bounds the loss-of-load probability. For planning, Chen et al. [chen2024multistage] proposed a method to do multi-stage planning using a chance constraint for unmet demand.
@@
Tractable Probabilistic Models. Probabilistic models have long provided a principled foundation for reasoning under uncertainty. These models describe a probability distribution over observed and latent variables, allowing one to answer questions such as: What is the likelihood of a given event? What is the expected value of a function under uncertainty? Or what is the most likely explanation for the observed data? This ability to represent and manipulate uncertainty is essential for decision-making in domains ranging from natural language processing and computer vision to power systems and finance. In fact, probability theory remains the most widely adopted framework for decision-making under uncertainty in artificial intelligence (AI) and machine learning (ML).
@@
Over the last two decades, the field has witnessed remarkable progress in developing expressive generative models. Deep generative approaches such as generative adversarial networks [goodfellow2014gans], variational autoencoders [kingma2022autoencodingvariationalbayes], and normalizing flows [papamakarios2021normalizingflowsprobabilisticmodeling] enabled closer to real-world data density estimation. These models show outstanding performance at generating highly realistic samples and capturing intricate dependencies in data, offering unprecedented modeling flexibility. Additionally, probabilistic programming languages and advances in Bayesian nonparametrics have expanded the expressiveness of probabilistic modeling frameworks.
@@
However, the increase in expressiveness comes at a significant computational cost. Inference in these models (computing marginals, conditionals, expectations, or the most likely assignment of variables) is typically intractable. For most deep generative models, it is unfeasible to compute efficiently the probability of a given outcome; instead, inference relies on sampling or approximate optimization. These procedures can be slow, unstable, or provide no guarantees. As [choi2020probabilistic] highlights, this creates a paradox: while models become better at fitting data distributions with high fidelity, they become less reliable for reasoning about those distributions, undermining the very goal of probabilistic modeling, which is not only to mimic data but to support decision-making through accurate probability computations.
@@
Tractable probabilistic models are an alternative to deal with the abovementioned paradox. These are models designed explicitly so that inference queries can be computed exactly and in polynomial time. The idea is not new: there are some restricted forms of TPMs proposed a few decades ago, such as the Kalman filter [Kalman1960], hidden Markov models (HMMs) [Baum1966], and tree-structured distributions [chow1968dependencetrees]. Although these models had limited expressive power, they were highly influential in signal processing, speech recognition, and computational biology thanks to their ability to support efficient inference algorithms.
@@
TPM vs. Other Probabilistic Models. Compared to mixture models (e.g., Gaussian mixtures), TPMs can scale to higher dimensions without exponential blow-up in the number of mixture components. Compared to deep generative models, TPMs provide sampling and exact probability evaluation, which is critical in settings requiring probabilistic guarantees—such as safety verification in autonomous systems, reliability analysis in power grids, or enforcing chance constraints in stochastic optimization. Monte Carlo simulation can approximate such probabilities, but may require enormous sample sizes; TPMs can compute them directly and precisely from the model representation.
TPMs in Chance-Constrained Optimization. TPMs allow efficient evaluation of probabilities, making them ideal candidates for implementing chance constraints in optimization. A notable recent example is Němeček et al. [nemecek2024generating], who encode an SPN into a mixed-integer optimization formulation (LiCE), enabling probability-based constraints on counterfactual likelihoods. Additionally, Ghandi et al. [ghandi2024probabilistic] propose methods to integrate propositional constraints directly into probabilistic circuits using convex optimization, refining the circuit distribution to satisfy logical constraints without retraining. Together, these works provide a foundation for embedding probabilistic constraints—such as enforcing reliability or safety with high probability—inside optimization models, suggesting TPMs could support chance-constrained expansion planning in energy systems.
@@ 3.1 Chance-Constrained Capacity/Generation/Storage Expansion Problems
Chance-Constrained Capacity/Generation/Storage Expansion Problems with TPMs. The main challenges that emerge from the problem described in (4) are two-fold: 1) we need a reliable model for the chance constraint, and 2) we need to be able to perform inference over it many times. Fortunately, TPMs offer a feasible alternative to tackle both challenges, first, by allowing for efficient estimation of the chance constraint, and second, by allowing the integration into optimization pipelines as shown in the work of [nemecek2024generating]. To enable the integration in such a framework, we start by defining a binary random variable $Y=\mathds{1}(G(x,\xi)\geq 0)$ which allows expressing the chance constraint as follows:
@@
Optimizing TPM-Based Chance Constraint. In the energy industry, most large-scale planning and operational problems are formulated as linear or mixed-integer linear programs (LP/MILP) because of their computational efficiency and compatibility with commercial solvers. These formulations dominate applications such as generation and transmission expansion, unit commitment, and market clearing [conejo2010decision, kirschen2018fundamentals]. Nonlinear formulations—such as AC optimal power flow—are typically linearized to maintain convexity and guarantee scalability. More recent work has extended these methods through stochastic and robust optimization frameworks [chenRobustOptimizationTransmission2014, roldanRobustTransmissionNetwork2019], enabling planners to explicitly account for uncertainty while preserving the tractable structure required for practical implementation. It is natural to require that our proposed chance-constrained approach be compatible with LP/MILP solvers. As shown in the work of Němecěk et al. [nemecek2024generating], we can achieve this by picking a TMP that could translate to linear constraints, in their case, an SPN, which is our selected approach as well. An SPN is a structured probabilistic model that represents a complex joint distribution through a combination of weighted sums and products of simpler distributions.
@@
These structural constraints for sum nodes ensure that SPNs support exact marginalization and conditioning in time linear regarding the network size. As a result, SPNs can efficiently compute densities and expectations required in stochastic and chance-constrained optimization problems. However, to enable the use of an SPN in the LP/MILP setup, we need to encode them into linear constraints, which is non-trivial, and some additional considerations are required. Following the approach of Němeček et al. [nemecek2024generating], changing all computations in logarithmic space simplifies the formulation of product nodes and improves numerical stability.
@@
This relation is purely linear and transfers directly into linear programming.
@@
In this formulation, the approximation provides a lower bound of the actual log-sum-exp value, sufficient when maximizing likelihood or probability mass.
@@
In contrast, the proposed TPM-based formulation decouples the optimization problem from the explicit enumeration of scenarios. With a trained SPN, evaluating or marginalizing probabilities required for the chance constraint involves only a fixed number of inference queries, each computable in time linear with respect to the network size but independent of $K$ and effectively replacing scenario enumeration with probabilistic inference, enabling the optimizer to reason over the full uncertainty distribution without increasing the model’s dimensionality. Moreover, because TPMs train on arbitrary numbers of samples, their accuracy improves with larger datasets—thus, unlike scenario-based methods, increasing $K$ enhances model quality rather than computational cost.

### v2

@@ Abstract
This framework enables direct embedding of chance-constrained optimization into investment planning, enforcing safety or reliability with prescribed confidence levels. TPMs allow both scenario analysis and volatility quantification by compactly representing high-dimensional uncertainties. We demonstrate the effectiveness of the approach through a representative power system planning case study, illustrating its computational and reliability advantages over traditional scenario-based models.
@@ 1 Introduction
With the increasing and more tangible effects of climate change, policymakers are intensifying demands for cleaner power systems, compelling grid operators to accelerate the integration of renewable energy into their networks. While renewables provide a feasible pathway to decarbonization and meeting greenhouse gas reduction targets, their weather-dependent and stochastic nature introduces profound operational challenges. Reliability is paramount: supply interruptions can propagate through the economy, disrupt critical infrastructure, and undermine social welfare. At the same time, the variability of renewable sources increases the need for additional reserves, flexibility options, and grid reinforcements, all of which carry high costs that grow with penetration levels [edenhofer2011renewable, impram2020challenges]. To maintain cost-effectiveness and resilience, system operators must ensure electricity remains affordable while safeguarding against blackouts and extreme events. These dual requirements have made rigorous planning indispensable, with infrastructure investments carefully evaluated to minimize the risk of decisions that could compromise long-term security and resilience of supply.
@@
Planning in power systems has long focused on determining the amount of new capacity to build, when to deploy it, and which technologies are most suitable to meet projected demand growth. Traditionally, modelling these generation and transmission expansion problems relies on deterministic assumptions, where long-term forecasts of demand, fuel prices, and technology costs are fixed inputs [taylor2015convex, kirschen2018fundamentals]. Over time, however, it became evident that such assumptions overlook the fundamental uncertainty associated with economic growth, policy developments, the impacts of climate change, and demand response [paragElectricityMarketDesign2016a]. Consequently, more recent approaches emphasize the direct incorporation of uncertainty into planning frameworks to ensure that expansion strategies remain viable under a wide range of possible futures.
@@
To address these challenges, planning methodologies have evolved towards stochastic programming and robust optimization [ben2009robust, conejo2010decision, pflug2014multistage, hadjidimitriou2021mathematical], which provide systematic ways of evaluating investment decisions under uncertain demand growth and variable renewable generation [zakaria2020uncertainty, 10931726]. In parallel, structural changes in the sector are shifting attention from large-scale, centralized capacity additions toward the integration of distributed energy resources (DERs), such as rooftop solar, storage systems, and demand response, which require more decentralized and flexible planning approaches [wang2013stochastic, asensio2016joint, MOTTA2024401]. Together, these developments reflect a broader shift in power system planning, where the traditional focus on least-cost expansion is now complemented by the need to ensure system resilience and long-term reliability in the face of uncertain and rapidly changing operating conditions.
@@
In this work, we address the challenges of long-horizon investment planning in power utilities, such as generation and transmission expansion planning, where operators must make decisions decades in advance under profound uncertainty. Forecasting the future energy mix and consumption patterns over such horizons is inherently complex, and current practice often relies on a limited set of discrete scenarios that represent key trajectories of exogenous variables. Statistically, this corresponds to a mixture of Dirac measures, which allows planners to capture a handful of plausible futures but offers little insight into the variability and volatility within each scenario. We propose replacing this paradigm with inference over TPMs integrated into the optimization problem. By compactly representing distributions over high-dimensional uncertainties, TPMs enable reasoning about both scenarios and the associated volatility, and the inference is representable in mixed-integer optimization. We illustrate this approach with a representative case study, demonstrating its feasibility for expansion planning under uncertainty and highlighting its potential for broader application across diverse planning contexts in the energy sector.
@@ 2 Background
Planning in Power Systems. Planning is a cornerstone of power system operation and development, ensuring that electricity supply can meet future demand reliably and at reasonable cost. Because power infrastructure has long lifetimes and requires significant capital investment, planning decisions must be made years or even decades in advance, often under deep uncertainty about technology, policy, and demand evolution. This task has traditionally been formulated as a capacity expansion problem, typically divided into generation expansion planning [DeJonghe2012830, Oderinwale2020] (GEP), transmission expansion planning [Xie2020, Lschenbrand2021] (TEP), and distribution expansion planning [Anjo2018, Gbadamosi2020] (DEP). Each has its own time horizons and decision variables. Collectively, these planning activities form the backbone of long-term power system development, linking near-term operational security with long-term goals such as decarbonization, resilience, and affordability [MOTTA2024401]. Traditional formulations often adopt deterministic structures that model uncertainty under a single problem, containing a set of scenarios, which are limited by computational power, leading to the common practice of selecting the most representative cases. As long-term planning increases uncertainty, it is typical to consider certain factors, such as focusing solely on one of the expansion problems, disregarding unit commitment constraints, or modeling the transmission system using a DC Optimal Power Flow (DCOPF) approach. The DCOPF is a linearized approximation of the AC power flow equations that simplifies the AC power flow equations, enabling scalable optimization for large planning problems, albeit at the expense of detailed network accuracy [doi:10.1049/iet-gtd.2015.0561]. More recent work emphasizes the need to explicitly capture uncertainty in demand and external drivers such as climate change. For example, the work in [LI2016401] models discrete climate scenarios within a stochastic generation expansion framework to evaluate viable investment strategies across multiple futures.
@@
Together, $X$ and $\Xi$ incorporate uncertainty into the investment decision process to make a final decision that aligns with the operator’s primary objective: minimizing costs while maintaining the system’s reliability. However, as the space of possible values for each $\xi_{t}$ is extremely large, it is prohibitive to explore them all; thus, it is common to build an estimate ($\hat{\Xi}$) with some representative cases and optimize $x$ over it, yielding to solving one constrained optimization problem for each $\xi\sim\hat{\Xi}$ as follows:
@@
to posteriorly pick the least expensive solution among the selected scenarios. The constraints in (3) represent the various considerations that limit the minimization of the primary objective (investment cost) while ensuring compliance with all scenario-dependent operational and policy constraints. This compact notation allows the formulation to encompass various system constraints without explicitly enumerating each one.
@@
Relevant Work on Scenario Expansion Planning. In power system planning, stochastic optimization and robust methods represent two complementary pathways to deal with uncertainty. For instance, Giannelos et al. [giannelos2025stochastic] propose a multi-stage stochastic network expansion model that accounts for both exogenous and endogenous uncertainties, capturing how investments evolve over time under varying conditions. On the robust side, Chen et al. [chenRobustOptimizationTransmission2014] introduced minimax cost and regret metrics to guard against worst-case outcomes. In contrast, Roldán et al. [roldanRobustTransmissionNetwork2019] extended that line by modelling correlated uncertainty sets, making the robust expansion plans more realistic. Other works further refine robustness: Li et al. [Li2016] incorporate ramping uncertainty and construction delays, and Velloso et al. [velloso2018distributionally] move toward distributionally robust formulations that sit between full stochastic and worst-case models. More recently, Brenner et al. [brenner2024learning] have combined stochastic planning with learning-based approximations to address computational complexity in large-scale systems. In parallel, Gholami et al. [gholami2022dynamic] explore joint expansion of power and gas infrastructures under uncertainty, illustrating how interdependent systems can be co-optimized.
@@
Chance Constraint Optimization. A central difficulty in power system planning is that these decisions happen under uncertainty. Traditional deterministic formulations provided valuable insights but offered limited protection against the wide range of plausible futures. As the penetration of renewables grows and climate-related risks become more pronounced [paragElectricityMarketDesign2016a], planning frameworks increasingly require explicit methods to incorporate uncertainty and evaluate the robustness of proposed investment strategies. One promising direction is to include chance constraints, which involve enforcing constraints probabilistically rather than for all possible outcomes, an idea previously explored in operation optimization problems.
@@
Chance-constrained formulations have already existed to enhance short-term scheduling. For example, Tan et al. [tan2020dual] developed a dual-timescale generation scheduling framework that integrates wind power, demand response, and energy storage, using chance constraints to procure sufficient ramping capacity under uncertainty. Similarly, Wang et al [wang2013stochastic] studied stochastic unit commitment with uncertain demand response, where a random variable models customer price elasticity and a chance constraint bounds the loss-of-load probability. For planning, Chen et al. [chen2024multistage] proposed a method for multi-stage planning that utilizes a chance constraint to address unmet demand.
@@
Tractable Probabilistic Models. Probabilistic models have long provided a principled foundation for reasoning under uncertainty. These models describe a probability distribution over observed and latent variables, enabling one to answer questions such as: What is the likelihood of a given event occurring? What is the expected value of a function under uncertainty? Or what is the most likely explanation for the observed data? This ability to represent and manipulate uncertainty is essential for decision-making in domains ranging from natural language processing and computer vision to power systems and finance. In fact, probability theory remains the most widely adopted framework for decision-making under uncertainty in artificial intelligence (AI) and machine learning (ML).
@@
Over the last two decades, the field has witnessed remarkable progress in developing expressive generative models. Deep generative approaches, such as generative adversarial networks [goodfellow2014gans], variational autoencoders [kingma2022autoencodingvariationalbayes], and normalizing flows [papamakarios2021normalizingflowsprobabilisticmodeling], enable closer approximation to real-world data density estimation. These models show outstanding performance at generating highly realistic samples and capturing intricate dependencies in data, offering unprecedented modeling flexibility. Additionally, probabilistic programming languages and advances in Bayesian nonparametrics have expanded the expressiveness of probabilistic modeling frameworks.
@@
However, the increase in expressiveness comes at a high computational cost. Inference in these models (computing marginals, conditionals, expectations, or the most likely assignment of variables) is typically intractable. For most deep generative models, it is unfeasible to compute efficiently the probability of a given outcome; instead, inference relies on sampling or approximate optimization. These procedures can be slow, unstable, or provide no guarantees. As [choi2020probabilistic] highlights, this creates a paradox: while models become better at fitting data distributions with high fidelity, they become less reliable for reasoning about those distributions, undermining the very goal of probabilistic modeling, which is not only to mimic data but to support decision-making through accurate probability computations.
@@
Tractable probabilistic models provide an alternative approach to addressing the aforementioned paradox. These models are designed explicitly to enable the exact computation of inference queries in polynomial time. The idea is not new: restricted forms of TPMs were proposed a few decades ago, such as the Kalman filter [Kalman1960], hidden Markov models (HMMs) [Baum1966], and tree-structured distributions [chow1968dependencetrees]. Although these models had limited expressive power, they were highly influential in signal processing, speech recognition, and computational biology thanks to their ability to support efficient inference algorithms.
@@
TPM vs. Other Probabilistic Models. Compared to mixture models (e.g., Gaussian mixtures), TPMs can scale to higher dimensions without exponential blow-up in the number of mixture components. Compared to deep generative models, TPMs offer sampling and exact probability evaluation, which is crucial in settings that require probabilistic guarantees—such as safety verification in autonomous systems, reliability analysis in power grids, or enforcing chance constraints in stochastic optimization. Monte Carlo simulation can approximate such probabilities, but it may require enormous sample sizes. In contrast, TPMs can compute them directly and precisely from the model representation.
TPMs in Chance-Constrained Optimization. TPMs allow efficient evaluation of probabilities, making them ideal candidates for implementing chance constraints in optimization. A notable recent example is LiCE [nemecek2024generating], which encodes an SPN into a mixed-integer optimization formulation, enabling constraints on the likelihoods of counterfactual explanations of AI models. Additionally, Ghandi et al. [ghandi2024probabilistic] propose methods to integrate propositional constraints directly into probabilistic circuits using convex optimization, refining the circuit distribution to satisfy logical constraints without retraining. Together, these works provide a foundation for embedding probabilistic constraints—such as enforcing reliability or safety with high probability—inside optimization models, suggesting TPMs could support chance-constrained expansion planning in energy systems.
@@ 3.1 Chance-Constrained Capacity/Generation/Storage Expansion Problems
Chance-Constrained Capacity/Generation/Storage Expansion Problems with TPMs. The main challenges that emerge from the problem described in (4) are two-fold: 1) we need a reliable model for the chance constraint, and 2) we need to be able to perform inference over it many times. Fortunately, TPMs offer a feasible alternative to tackle both challenges, first, by allowing for the efficient estimation of the chance constraint, and second, by enabling integration into optimization pipelines, as shown in the work of [nemecek2024generating]. To enable the integration in such a framework, we start by defining a binary random variable $Y=\mathds{1}(G(x,\xi)\geq 0)$ which allows expressing the chance constraint as follows:
@@
Optimizing TPM-Based Chance Constraint. In the energy industry, most large-scale planning and operational problems are formulated as linear or mixed-integer linear programs (LP/MILP) because of their computational efficiency and compatibility with commercial solvers. These formulations dominate applications such as generation and transmission expansion, unit commitment, and market clearing [conejo2010decision, kirschen2018fundamentals]. Nonlinear formulations—such as AC optimal power flow—are typically linearized to maintain convexity and guarantee scalability. More recent work has extended these methods through stochastic and robust optimization frameworks [chenRobustOptimizationTransmission2014, roldanRobustTransmissionNetwork2019], enabling planners to explicitly account for uncertainty while preserving the tractable structure required for practical implementation. It is natural to require that our proposed chance-constrained approach be compatible with LP/MILP solvers. As shown in the work of Němecěk et al. [nemecek2024generating], we can achieve this by selecting a TMP that can be translated into linear constraints, in their case, an SPN, which is our chosen approach as well. An SPN is a structured probabilistic model that represents a complex joint distribution through a combination of weighted sums and products of simpler distributions.
@@
These structural constraints for sum nodes ensure that SPNs support exact marginalization and conditioning in linear time with respect to the network size. As a result, SPNs can efficiently compute densities and expectations required in stochastic and chance-constrained optimization problems. However, to enable the use of an SPN in the LP/MILP setup, we need to encode them into linear constraints, which is non-trivial, and some additional considerations are required. Following the approach of Němeček et al. [nemecek2024generating], changing all computations in logarithmic space simplifies the formulation of product nodes and improves numerical stability.
@@
This relationship is purely linear and can be directly applied to linear programming.
@@
In this formulation, the approximation provides a lower bound on the actual log-sum-exp value, which is sufficient when maximizing likelihood or probability mass.
@@
In contrast, the proposed TPM-based formulation decouples the optimization problem from the explicit enumeration of scenarios. With a trained SPN, evaluating or marginalizing probabilities required for the chance constraint involves only a fixed number of inference queries, each computable in time linear with respect to the network size but independent of $K$ and effectively replacing scenario enumeration with probabilistic inference, enabling the optimizer to reason over the full uncertainty distribution without increasing the model’s dimensionality. Moreover, because TPMs train on arbitrary numbers of samples, their accuracy improves with larger datasets; thus, unlike scenario-based methods, increasing $K$ enhances model quality rather than incurring additional computational cost.